OpenStudy (anonymous):

If f(x) = 4x – 6 and g(x) = 2x, find

OpenStudy (shamil98):

(4x - 6)(2x) just distribute the 2x .

OpenStudy (anonymous):

f(x) = 4x – 6 and g(x) = 2x Therefore, \[f(x) . g(x) =( 4x – 6) 2x = 4x \times 2x - 6 \times 2x = 8x^2- 12x\] @sirizarry78

OpenStudy (anonymous):

thank you @dpasingh